# DataPlatter: Boosting Robotic Manipulation Generalization with Minimal Costly Data

### v1

@@ Abstract
The growing adoption of Vision-Language-Action (VLA) models in embodied AI intensifies the demand for diverse manipulation demonstrations. However, high costs associated with data collection often result in insufficient data coverage across all scenarios, which limits the performance of the models. It is observed that the spatial reasoning phase (SRP) in large workspace dominates the failure cases. Fortunately, this data can be collected with low cost, underscoring the potential of leveraging inexpensive data to improve model performance. In this paper, we introduce the DataPlatter method, a framework that decouples training trajectories into distinct task stages and leverages abundant easily collectible SRP data to enhance VLA model’s generalization. Through analysis we demonstrate that sub-task-specific training with additional SRP data with proper proportion can act as a performance catalyst for robot manipulation, maximizing the utilization of costly physical interaction phase (PIP) data. Experiments show that through introducing large proportion of cost-effective SRP trajectories into a limited set of PIP data, we can achieve a maximum improvement of 41% on success rate in zero-shot scenes, while with the ability to transfer manipulation skill to novel targets.
footnotetext: 1 Meituan Inc.footnotetext: 2 Institute of Computing Technology, Chinese Academy of Sciences.footnotetext: Corresponding authors.
@@ 1 Introduction
Our key insight stems from two critical observations: (1) The spatial understanding ability required in SRP exhibits higher environmental variability compared to PIP, since the manipulation stage for a specified target is relatively fixed with little correlation with the surrounding scene; (2) Neural networks demonstrate distinct attention patterns during different task stages, such as the focus on target’s location and spatial occupancy to avoid collision in SRP while shifted to the target proportion in PIP. These findings suggest that sub-task-specific training strategies could better align with the model’s learning characteristics, utilizing varying proportions of these data segments across the sub-tasks.
Both Tan et al. [41] and our experiments (see Tab. 3) have demonstrated that a smaller workspace can significantly improve the success rate of operating tasks. This indicates that decoupling operation stages with different centers of attention can improve generalization performance. Furthermore, this variation in sub-task difficulty can cause the model to overfit on the simpler, small-workspace stage while underfit on the large-workspace stage, which necessitates different data volumes for each stage. In this paper, we propose the DataPlatter method, which decouples training data across different operation stages, constructs a implicit sub-task specific training procedure, and leverages a large amount of easily collectible SRP data to train this stage, to improve the performance of VLA models. With our method, a substantial amount of labor-intensive teleoperation time traditionally required to collect complex manipulation trajectories, e.g. the 17 months needed for 130k episodes in RT-1 [3], can be significantly reduced. Instead, program-driven automatic collection can be employed to acquire a large volume of low-interaction trajectories in extensive workspaces. This approach not only reduces manual efforts but also greatly enhances the potential to leverage larger datasets to improve model capabilities.
@@
  • We introduced the DataPlatter methods, which utilizes additional cost-effective SRP trajectories to improve the model’s generalization performance in zero-shot scenes.
@@
  • Decoupling task stages in a dataset to build implicit sub-task specific training processes can offer a flexible approach to enhancing model performance at the sub-task level.
@@ 2 Related Work
Robot manipulation datasets
@@
Currently, most EAI models are limited to executing tasks they have explicitly encountered during training. For instance, even if a model be trained to pick up bottles, it cannot generalize this to pick up a cola can. Although this problem have already been studied through methods ranging from early domain randomization [13], meta-learning [8] and data augmentation[17] to recent advancements in world model building [5, 27] and spatial reasoning [25, 12], the generalization performance on out-of-distribution (OOD) novel targets still shows limited improvement. [5, 53] try to transfer the world knowledge from large models trained with Internet-scale data to robot action reasoning, but the the manipulation experience of OOD targets from “practicing” can not be efficiently acquired from “reading”, while [27, 38, 38] are trying to directly use the general ability to guide the agent’s action logic. [25, 12, 51] are working improving the action performance through understanding the spatial information in the workspace. Zhu et al. [52] transfer the target knowledge to similar objects through text-image pairs, but still needs auxiliary information to get a better performance during inference. In this paper we propose an end-to-end training method, which can improve the generalization performance on OOD targets by a large margin.
3 Method
@@
As illustrated in Fig. 2, this paper introduces the DataPlatter method, which segments robot manipulation trajectories into spatial reasoning and physical interaction phases according to the extent of the agent’s interaction with objects in the environment. By employing a mixture of two-stage data in appropriate proportions, we aim to achieve a generalization performance comparable to using complete data for model training. This approach effectively reduces the reliance on expensive PIP data.
@@
3.1 Problem Formulation
In this paper we focus on the VLA models that utilize behavior cloning, which is a category of IL methods. Consider a robot manipulation trajectory dataset $\mathcal{D}^{F}=\left\{\tau_{i}^{F}\right\}_{i=1}^{N}$, where each full-stage trajectory $\tau_{i}^{F}=\left\{l^{i},o_{1}^{i},a_{1}^{i},o_{2}^{i},\cdots,a_{T-1}^{i},o_{T}^{i}\right\}$ consist of the task’s language instruction $l$, the agent’s observation $o_{t}^{i}$ at each time step $t$, and the action $a_{t}^{i}$ taken by the agent. The VLA model $\Psi_{\theta}$ with parameter $\theta$ takes as input the task instruction and a segment of observation history $\mathcal{O}_{t,L}^{i}=\left\{o_{t-L+1},\cdots,o_{t}\right\}$ of length $L$, and predicts the action chunk $\mathcal{A}_{t,L,H}^{i}=\left\{a_{t-L+1},\cdots,a_{t+H}\right\}$ that the agent should execute to accomplish the task in the past $L$ and next $H$ time steps, i.e.
$\hat{\mathcal{A}}_{t,L,H}^{i}$ $=\Psi_{\theta}\left(\mathcal{O}_{t,L}^{i},l^{i}\right)$ (1)
$=Dec\left(LLM\left(Enc_{v}\left(\mathcal{O}_{t,L}^{i}\right),Enc_{l}\left(l^{i}\right)\right)\right),$
where $Enc_{v}$ and $Enc_{l}$ are vision and language encoders, $LLM$ is the pretrained large language model, and $Dec$ donates the action decoder. Typically, vision encoders like CLIP [36] are pretrained using image-text pairs to provide aligned visual-textural semantics, facilitating seamless integration with LLMs, and are generally kept frozen during training of the VLA model. The LLM such as GPT [1] or LLaMA [32] serves as the core of the model due to its robust general reasoning capabilities and typically employs adapters [20, 26] for integrating multi-modal input tokens. Action decoders usually consist of several lightweight neural layers that interpret the action token chunks output by the LLM and transform them into physically meaningful actions, e.g. 6-DoF poses of end effectors.
The objective of model optimization is to minimize the discrepancy between predicted and demonstration action sequences, i.e.
$\theta^{*}=\arg\min_{\theta}Err\left(\hat{\mathcal{A}}_{t,L,H}^{i},\mathcal{A}_{t,L,H}^{i}\right),$ (2)
where $Err\left(\cdot\right)$ is a function that measures the divergence between the demonstrated and predicted actions, and has different representations for different action forms.
3.2 VLA Training with DataPlatter
To utilize the trajectories of different sub-tasks in the dataset, we first segment the given full-stage trajectory $\tau_{i}\in\mathcal{D}$ into SRP and PIP based on the distance between the end-effector $G$ and the target object $T$, as well as the visibility of the target in the wrist camera $C_{w}$. It is assumed that there is one wrist camera and one static camera in the scene, which is a common configuration in most datasets. More formally, for a target object $T$ at position $p_{T}$, an end-effector $G$ at position $p_{G}$ and a wrist camera $C_{w}$ at pose $P_{C}=\left(p_{C},R_{C}\right)$ defined under OpenCV frame, where $p_{T},p_{G},p_{C}\in\mathbb{R}^{3}$ and $R_{C}\in SO(3)$, the PIP begins if
$\left\{\begin{matrix}\left\|p_{G}-p_{T}\right\|\leq d_{th},\\ \frac{v_{z}^{T}\cdot\left(p_{T}-p_{C}\right)}{\left\|p_{T}-p_{C}\right\|}>\arccos\frac{\alpha_{fov}}{2},\end{matrix}\right.$ (3)
where $v_{z}\in\mathbb{R}^{3}$ s.t. $\left\|v_{z}\right\|=1$ is the direction vector of $z$-axis of frame $P_{C}$, $d_{th}$ is the distance thresholds and $\alpha_{fov}$ is the field of view of $C_{w}$. The PIP stops once the interaction-rich manipulation stage is accomplished, e.g. after grasping the target in pick-and-place task or triggering the button in switch-operation task. Apart from the PIP, the rest of the trajectory is designated as the SRP. Following such procedure, the trajectory can be divided into several segments $\tau_{i}^{F}=\left\{\tau_{i,1}^{SRP},\tau_{i,1}^{PIP},\tau_{i,2}^{SRP},\cdots\right\}$. Correspondingly, the dataset can be divided into two sub-datasets: $\mathcal{D}^{F}=\mathcal{D}^{SRP}\cup\mathcal{D}^{PIP}$, where $\mathcal{D}^{SRP}=\left\{\tau_{i,j}^{SRP}\right\}$ contains all of the SRP segments in the trajectories and $\mathcal{D}^{PIP}=\left\{\tau_{i,j}^{PIP}\right\}$ contains the manipulation segments. Note that we are aiming to train the VLA model with a larger quantity of easily-collectible SRP data than expensive PIP data, so in practice independently-collected SRP dataset $\mathcal{D}_{ind}^{SRP}$ can be included in training.
@@ 3.2 VLA Training with DataPlatter
Before the training phase of the VLA model, we sample $N_{1}$ and $N_{2}$ segments in $\mathcal{D}$ and $\mathcal{D}_{ind}^{SRP}$ respectively, and construct a new dataset $\mathcal{D}^{Mix}$ to train the model, which in this paper we call DataPlatter, i.e.
@@ 4 Experiments
In this section, we investigate how the total number of training trajectories and the proportion of SRP data impact task success rates. We aim to identify the optimal strategy for leveraging easily accessible SRP data to enhance the generalization performance of the VLA model.
@@ 4.1 Environment Setup
VLA model
Training data
In the simulation environment of IsaacSim, we generated a dataset for object-picking tasks involving target objects of various categories and geometrical shapes. For SRP-only trajectories, to provide a implementable pipeline in real-world robots, we did not directly read the object information from simulation, instead we applied a detection-sampling method provided in Fig. 1(c). Details are shown in Sec. A.1. During trajectory generation, we observe that the SRP-only trajectories are generated 2.5×2.5$\times$2.5 × faster than those using full-stage data, while the length of the full-stage data is only 1.4×1.4$\times$1.4 × that of the SRP data. In real-world data collection this discrepancy can only be even larger. Other datasets we used in the experiments are divided using method provided in Eq. 3 with $d_{th}=0.2m$ and $\alpha_{fov}=\frac{\pi}{3}$, details are shown in Sec. A.2.
Evaluation
Training
@@
4.2 Experiment Results
@@
4.2.1 Generalization with SRP Data
Firstly, we aim to determine whether increasing the amount of SRP data can improve the model’s generalization performance. We randomly select the PIP trajectories of $M$ target objects as our task, and create a multi-task dataset to train the RoboMM models. For each task, we use a fixed number of $N_{1}=100$ full-stage trajectories $\tau_{i}^{F}\in\mathcal{D}^{F}$, alongside a variable number of $N_{2}$ SRP trajectories $\tau_{i}^{SRP}\in\mathcal{D}_{ind}^{SRP}$. In this experiment, we set $M\in\left\{5,20\right\}$ and $N_{2}\in\left\{0,100,200,300\right\}$. The results are presented in Fig. 3.
It can be observed that incorporating additional automatically collected SRP data during training significantly elevates the model’s success rate, achieving a maximum margin of 41% over the reference model. Meanwhile, incorporating SRP data can achieve an equivalent performance to manually collected full-stage data (marked with gray arrow) at substantially lower cost. However, the performance bottleneck occurs earlier than that of using full-stage data, when $N_{2}>2N_{1}$, i.e. in this experiment $N_{2}>200$, increasing the number of SRP trajectories $N_{2}$ while keeping $N_{1}$ constant yields no significant improvement in the model’s performance. This suggests that incomplete trajectories cannot be added indefinitely, as insufficient operational data may hinder the model from learning effective action logic, and the proportion of data mixture is discussed in Sec. 4.3.2. However, at this point, the generalization performance of a model trained with $N_{1}=300$ full-stage trajectories across 5 tasks can still improve by 7% ($68\%\rightarrow 75\%$) with an additional $N_{2}=300$ SRP trajectories. This performance is close to the bottleneck of 80% achievable with more full-stage data.

### v2

# Boosting Robotic Manipulation Generalization with Minimal Costly Data

@@ Abstract
The growing adoption of Vision-Language-Action (VLA) models in embodied AI intensifies the demand for diverse manipulation demonstrations. However, high costs associated with data collection often result in insufficient data coverage across all scenarios, which limits the performance of the models. It is observed that the spatial reasoning phase (SRP) in large workspace dominates the failure cases. Fortunately, this data can be collected with low cost, underscoring the potential of leveraging inexpensive data to improve model performance. In this paper, we introduce the RoboTron-Craft, a stage-divided and cost-effective pipeline for realistic manipulation generation. Base on this, the RoboTron-Platter method is introduced, a framework that decouples training trajectories into distinct task stages and leverages abundant easily collectible SRP data to enhance VLA model’s generalization. Through analysis we demonstrate that sub-task-specific training with additional SRP data with proper proportion can act as a performance catalyst for robot manipulation, maximizing the utilization of costly physical interaction phase (PIP) data. Experiments show that through introducing large proportion of cost-effective SRP trajectories into a limited set of PIP data, we can achieve a maximum improvement of 41% on success rate in zero-shot scenes, while with the ability to transfer manipulation skill to novel targets. Project available at https://github.com/notFoundThisPerson/RoboTron-Craft.
footnotetext: Meituan Inc.footnotetext: Corresponding authors.
@@ 1 Introduction
To answer these questions, we first introduce RoboTron-Craft, a stage-divided and cost-effective pipeline designed to generate physically realistic manipulation trajectories with photo-realistic observation, along with an environment for model evaluation. By utilizing this pipeline, we produce diverse trajectories across various stages, encompassing a wide range of objects and environments. Based on which, we reveals the logarithmic scaling law between VLA model performance and target object diversity, amount of training trajectories, as well as environmental diversity. Our key insight stems from two critical observations: (1) The spatial understanding ability required in SRP exhibits higher environmental variability compared to PIP, since the manipulation stage for a specified target is relatively fixed with little correlation with the surrounding scene; (2) Neural networks demonstrate distinct attention patterns during different task stages, such as the focus on target’s location and spatial occupancy to avoid collision in SRP while shifted to the target proportion in PIP. These findings suggest that sub-task-specific training strategies could better align with the model’s learning characteristics, utilizing varying proportions of these data segments across the sub-tasks.
Both Tan et al[41] and our experiments (see Tab. 4) have demonstrated that a smaller workspace can significantly improve the success rate of operating tasks. This indicates that decoupling operation stages with different centers of attention can improve generalization performance. Furthermore, this variation in sub-task difficulty can cause the model to overfit on the simpler, small-workspace stage while underfit on the large-workspace stage, which necessitates different data volumes for each stage. In this paper, we propose the RoboTron-Platter method, which decouples training data across different operation stages, constructs a implicit sub-task specific training procedure, and leverages a large amount of easily collectible SRP data to train this stage, to improve the performance of VLA models. With our method, a substantial amount of labor-intensive teleoperation time traditionally required to collect complex manipulation trajectories, e.g. the 17 months needed for 130k episodes in RT-1 [2], can be significantly reduced. Instead, program-driven automatic collection can be employed to acquire a large volume of low-interaction trajectories in extensive workspaces. This approach not only reduces manual efforts but also greatly enhances the potential to leverage larger datasets to improve model capabilities.
@@
  • We introduce the RoboTron-Craft, a stage-divided and cost-effective pipeline designed to generate realistic manipulation trajectories. Based on which we reveals the scaling law between model performance and trajectory diversity.
  • We introduced the RoboTron-Platter methods, which utilizes additional cost-effective SRP trajectories to improve the model’s generalization performance in zero-shot scenes.
@@ 2 Related Work
Robot datasets & benchmarks
@@
Currently, most EAI models are limited to executing tasks they have explicitly encountered during training. For instance, even if a model be trained to pick up bottles, it cannot generalize this to pick up a cola can. Although this problem have already been studied through methods ranging from early domain randomization [12], meta-learning [7] and data augmentation[16] to recent advancements in world model building [4, 24] and spatial reasoning [23, 11], the generalization performance on out-of-distribution (OOD) novel targets still shows limited improvement. [4, 53] try to transfer the world knowledge from large models trained with Internet-scale data to robot action reasoning, but the the manipulation experience of OOD targets from “practicing” can not be efficiently acquired from “reading”, while [24, 38, 38] are trying to directly use the general ability to guide the agent’s action logic. [23, 11, 51] are working improving the action performance through understanding the spatial information in the workspace. Zhu et al[52] transfer the target knowledge to similar objects through text-image pairs, but still needs auxiliary information to get a better performance during inference. In this paper we propose an end-to-end training method, which can improve the generalization performance on OOD targets by a large margin.
3 RoboTron-Craft Benchmark
To emphasize the ability of behavior cloning methods in manipulating various shaped objects in real world, a substantial number of demonstration trajectories is essential. Accordingly, as shown in Fig. 2, we introduce the RoboTron-Craft benchmark, which offers a cost-effective pipeline for generating physically realistic manipulation trajectories, along with an environment for model evaluation. Built upon the powerful NVIDIA Isaac Sim [34] platform, our benchmark can provide a smooth robot trajectory generation and evaluation pipeline on various commonplace objects, with photo-realistic rendering in multiple well-appointed scenes and reasonable physical feedback on interaction-rich tasks. Details are shown in Appendix A.
Figure 2: Overview of our data generation pipeline. (a) The environment is first initialized with multiple objects with random pose. (b) In the object-centric planning phase, feasible manipulation pose of the target is sampled. (c) In the agent-centric planning phase, our pipeline split the task into several segments, plan the execution trajectory and watch the task progress during execution.
3.1 Trajectory Generation Pipeline
We automatically generate robot demonstration trajectories on manipulation tasks with target objects in various geometry shapes, scattered on several workbenches in different rooms with various background textures. The generation pipeline is decomposed into a two-level hierarchy: object-centric planning and agent-centric planning, for the convenience of scalability across various task types.
Starting from randomly sampling a pre-defined environment and several manipulatable objects ${}\left\{\mathcal{M}\right\}$, our benchmark randomly places the objects in the scene with collision-free poses. In the following object-centric planning phase, our benchmark samples a target object ${}\mathcal{M}_{\text{tgt}}\in\left\{\mathcal{M}\right\}$, along with a feasible manipulation pose $\mathbf{g}=\left(\mathbf{R},\mathbf{t}\right)\in SE(3)$ of $\mathcal{M}_{\text{tgt}}$, which is sampled from a set of pose candidates $\mathcal{G}^{\text{cand}}_{\mathcal{M}}$ that are automatically generated using method described in Sec. A.4. Then during the agent-centric planning, task execution stages are supervised by the task manager with the known $\mathbf{g}$ and task template, and the collision-free robot trajectories in each stage are planned using motion planners with the priori knowledge about the geometry shape of the target and the spatial occupancy of the scene. Details can be found in Sec. A.2.
3.2 SRP Data Collection
Although it is efficient to collect demonstrations through our RoboTron-Craft pipeline in simulation, it is not implementable in real robot data collection, which is necessary to bridge the sim-to-real domain gap. On the other hand, during trajectory generation, time consumed by the motion planning when approaching the target and contact simulation when manipulating the target takes over 70% of the progress, eliminating these cost can greatly speed up data collection. Luckily, our experiment shows that enriching the spacial reasoning stage trajectories can promote scene generalization, meaning that we can collect the trajectory slice that only approaching the target without actual interaction.
To achieve this, starting from a ground truth manipulation pose sample $\mathbf{g}\in\mathcal{G}^{\text{cand}}_{\mathcal{M}}$ of target $\mathcal{M}_{\text{tgt}}$, an offset transform $\Delta\mathbf{T}\in SE(3)$, with random orientation offset $\Delta\mathbf{R}$ and translation offset $\Delta\mathbf{t}$ away from the object, is sampled and applied to $\mathbf{g}$, i.e.
$\mathbf{g}_{\text{app}}=\mathbf{g}\Delta\mathbf{T},$ (1)
to get the approaching pose $\mathbf{g}_{\text{app}}$ just before interacting with $\mathcal{M}_{\text{tgt}}$, which serves as the final state of our SRP trajectories. In this way, the need for planning the narrow path reaching exactly the target object for the motion planner as well as the complex collision computation of the physics simulator while interacting can be eliminated, which greatly saves the data collection time. Since the manipulation model can learn the correct interaction pose with the object, it will automatically fix the pose error while inference as long as the pose error is within a certain range, which is used as the sample space of $\Delta\mathbf{T}$. The parameters used in this stage are shown in Sec. A.5.
Further more, this procedure can be applied to the real-world robot trajectory collection. With sparse ground truth manipulation labels on the target object and an object pose estimation algorithm, SRP trajectory collection can be greatly accelerated since the robot can plan and execute the trajectories automatically, reducing the need for human tele-operation, thus the data richness can be greatly improved, providing the potential for better capability of VLA models in manipulation.
4 RoboTron-Platter: Training with Data Mixture
@@
As illustrated in Fig. 3, this paper introduces the RoboTron-Platter method, which segments robot manipulation trajectories into spatial reasoning and physical interaction phases according to the extent of the agent’s interaction with objects in the environment. By employing a mixture of two-stage data in appropriate proportions, we aim to achieve a generalization performance comparable to using complete data for model training. This approach can effectively reduce the reliance on expensive PIP data.
@@
In order to train the VLA model, we first need to collect a dataset of full-stage demonstration trajectories $\mathcal{D}^{F}=\left\{\tau^{F}_{i}\right\}_{i=1}^{N}$ for supervised training. Note that we are aiming to train the VLA model with a larger quantity of easily-collectible SRP data than expensive PIP data, so independently-collected SRP dataset $\mathcal{D}_{ind}^{SRP}$ can be included in our training process.
To utilize the trajectories of different sub-tasks in both $\mathcal{D}^{F}$ and $\mathcal{D}_{ind}^{SRP}$, we first segment the given full-stage trajectory $\tau_{i}^{F}\in\mathcal{D}^{F}$ into SRP and PIP based on the distance between the end-effector $G$ and the target object $T$. More formally, the PIP segment $\tau^{PIP}_{i}$ starts from the state that $G$ is at the approaching pose $\mathbf{g}_{\text{app}}$, and ends up when the interaction-rich stage is accomplished, e.g. after grasping the target in pick-and-place task or triggering the button in switch-operation task. The remaining segments of $\tau_{i}^{F}$ such as the approaching stage are all classified into SRP segments $\tau^{SRP}_{i}$. Following such procedure, the trajectory can be divided into several segments $\tau_{i}^{F}=\left\{\tau_{i,1}^{SRP},\tau_{i,1}^{PIP},\tau_{i,2}^{SRP},\cdots\right\}$. Correspondingly, the dataset can be divided into two sub-datasets: $\mathcal{D}^{F}=\mathcal{D}^{SRP}\cup\mathcal{D}^{PIP}$, where $\mathcal{D}^{SRP}=\left\{\tau_{i,j}^{SRP}\right\}$ and $\mathcal{D}^{PIP}=\left\{\tau_{i,j}^{PIP}\right\}$.
@@
Before the training phase of the VLA model, we sample $N_{1}$ and $N_{2}$ segments in $\mathcal{D}^{F}$ and $\mathcal{D}_{ind}^{SRP}$ respectively, and construct a new dataset $\mathcal{D}^{Mix}$ to train the model, which in this paper we call RoboTron-Platter, i.e.
@@ 5 Experiments
In this section, we investigate how the total number of training trajectories and the proportion of SRP data impact task success rates. We aim to identify the optimal strategy for leveraging easily accessible SRP data to enhance the generalization performance of the VLA model. Moreover, we investigate the scaling law of training data amount and the diversity of target geometries as well as environment setups.
@@ 5.1 Environment Setup
Unless otherwise stated, in this paper the RoboMM [47] is used as our baseline. Language instruction, RGB images from wrist and table cameras, and camera parameters are feed into the model in the training phase. During training we form the mixed dataset $\mathcal{D}^{Mix}$ with a varying of proportion of independent SRP segments, to verify the effectiveness of our RoboTron-Platter method, i.e.
@@
In this paper we use trajectories pick-and-place task on different target objects as our training data. The trained models are evaluated in our RoboTron-Craft benchmark using IsaacSim platform. A trail is considered succeeded if the agent successfully picked up the instruction-specified target object under the actions generated by the VLA model. Note that depth images are not utilized in evaluation, which is only used in training for supervision together with action trunks. More details are shown in Sec. C.1.
5.2 Experiment Results for RoboTron-Platter
5.2.1 Generalization with SRP Data
Firstly, we aim to determine whether increasing the amount of SRP data can improve the model’s generalization performance. We randomly select the PIP trajectories of $M$ target objects as our task, and create a multi-task dataset to train the RoboMM models. For each task, we use a fixed number of $N_{1}=100$ full-stage trajectories $\tau_{i}^{F}\in\mathcal{D}^{F}$, alongside a variable number of $N_{2}$ SRP trajectories $\tau_{i}^{SRP}\in\mathcal{D}_{ind}^{SRP}$. The results are presented in Fig. 4. Models trained with totally full-stage (FS) data ($N_{2}=0$) are also presented as a reference, to serve as the upper bound of model performance.
@@ 5.2.1 Generalization with SRP Data
[…]